OpenStudy (anonymous):

What do you do -n+7=12

OpenStudy (anonymous):

Do you want me to walk you through the steps to figure this out?

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

Gotta show work to

OpenStudy (anonymous):

Alright, first you subtract the 7 on both sides. So, -n + 7 - 7= 12 - 7

OpenStudy (anonymous):

I did that

OpenStudy (anonymous):

Confused what to divide by

OpenStudy (anonymous):

and then, your final answer will be -n = 5. You don't divide in this type of problem since it is only a single variable.

OpenStudy (anonymous):

So n=s just 5

OpenStudy (anonymous):

-n = 5. So yeah, that's just it.

OpenStudy (anonymous):

Here's another one m over 2 + 5 =s -3

OpenStudy (anonymous):

First, add the 2 and the 5 so you get 7= s - 3.

OpenStudy (anonymous):

And then you add the three on both sides, so 7 + 3 = s -3 + 3. And then you get 10 = s

OpenStudy (anonymous):

So add 5+2 an 3+2

OpenStudy (anonymous):

No, you add the the 5 and the 2 to get 7 and then your new equation is 7= s - 3

OpenStudy (anonymous):

So you do 5-5 since add isnopposite

OpenStudy (anonymous):

What u add both sides with?

OpenStudy (anonymous):

you add 3 to both sides, 7 + 3 = s - 3 + 3

OpenStudy (anonymous):

You're trying to get the variable by itself. So then you get 10 = s and that's it

OpenStudy (anonymous):

So add th my subtract

OpenStudy (anonymous):

I got 8 5+3 is 8

OpenStudy (anonymous):

U not suppose to subtract instead of add cuz I thought you was suppose to do the opposite of the sign

OpenStudy (anonymous):

I think you're a bit confused, which is okay. 2 + 5 = s - 3 7 = s - 3 add 3 to both sides: 7 + 3 = s - 3 + 3 10 = s

OpenStudy (anonymous):

So I added 5+2 which is 7 what is the next addition problem

OpenStudy (anonymous):

I'm suppose to marooned addition part of it out cuz it's suppose to equal 0 or be the same number than u just mark it out

OpenStudy (anonymous):

I got 5+2 down just confused on what else I'm suppose to add 2 with

OpenStudy (anonymous):

the 5 + 2 was part of the original equation so you don't add it to anything else you just simplify it so it equals 7

OpenStudy (anonymous):

Ok what r u suppose to divide it by?

OpenStudy (anonymous):

If u didn't simply it what would you have to put down next to the 5+2 what other problem

OpenStudy (anonymous):

There's no division required since it's just a single variable

OpenStudy (anonymous):

So m=s 7

OpenStudy (anonymous):

10 = s You want to find the value of s

OpenStudy (anonymous):

and s = 10

OpenStudy (anonymous):

There isn't a s on here tho loll hers no other letter but m

OpenStudy (anonymous):

2 + 5 =s -3 that was the original equation you asked.

OpenStudy (anonymous):

No m over two means divide + 5=a-3

OpenStudy (anonymous):

=-3 not a-3

OpenStudy (anonymous):

I'm not sure what you're talking about anymore, I think my work here is done.

OpenStudy (anonymous):

An I'm trying to find m

OpenStudy (anonymous):

Okay thanks

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."